Daniel is 69 years old and Ishaan is 6 years old.
How many years will it take until Daniel is only 4 times as old as Ishaan?

Answer by swincher4391(1107)

Let x be the number of years it takes for Daniel to be 4 times as old as Ishaan.

Then, 69+x = 4(6+x)
69+x = 24+4x
45 = 3x
x = 15
i.e. Daniel will be 84 and Ishaan will be 21.